OpenStudy (anonymous):

Phil has 11 stamps of denominations $0.37 and $0.23. If the total value of the stamps is $3.23 how many $0.37 stamps does Phil have?

OpenStudy (anonymous):

let phil has x number of stamps of $0.37 denomination then he will be having (11-x) number of stamps of @0.23 denomination in this case \[{{0.37}\times{x}}+{{0.23}\times(11-x)} = 3.23\] now solve this equation

OpenStudy (anonymous):

Is it -0.0371??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[{0.37x}-{0.23x}= 3.23 - 0.23\times11\] \[0.14x = 3.23 - 2.53\] \[0.14x = 0.7\] \[x= {{0.7}\over{0.14}}=5\]

OpenStudy (anonymous):

ohh I see!! Thanks ;))